A Php1,000, 9.5% MWDR Inc. bond has a semi-annual coupon rate of 15%. This matures after 6 years and the required rate of return is 8%. Find the bond price.

1
Expert's answer
2021-11-26T05:19:05-0500

bond price = "\sum\frac{C_n}{(1+YTM)^n}+\frac{P}{(1+i)^n}"

where

n is period which takes values from 0 to the nth period till the cash flows ending period,

Cn is coupon payment in the nth period,

YTM is interest rate or required yield,

P is par value of the bond.


"n=6\cdot2=12"

"C_n=1000\cdot0.15=Php\ 150"

"YTM=0.08\cdot6/12=0.04"

"P=1000\cdot0.095=Php\ 95"


bond price = "\displaystyle{\sum^{12}_{n=1}}\frac{150}{(1+0.04)^{12}}+\frac{95}{(1+0.15)^{12}}=Php\ 1585.31"